The following reaction was monitored as a function of time:
A→B+C
A plot of ln[A] versus time yields a straight line with slope −4.5×10−3 /s .

If the initial concentration of A is 0.240 M , what is the concentration after 240 s ?
Answer #1

For a first order reaction,a plot of ln[A] vs time gives a straight line with a negative slope and slope gives us the value of rate constant.

slope = -k

-4.5×10-3 s-1 = -k

4.5×10-3 s-1 = k

0.0045 s-1 = k

Also for the first order reaction,integrated rate law can be given by,

[A]t = [A]o e-kt

Where,

[A]t = concentration after t = 240 s

[A]o = initial concentration = 0.240 M

t = time = 240 s

[A]t = [A]o e-kt

[A]t = 0.240 × e-0.0045×240

[A]t = 0.240 × 0.339

[A]t = 0.082 M